1. Find the vector form of the general solution of the given linear system Ax = b ; then use that result to find the vector form of the general solution of Ax = 0.

x subscript 1 + x subscript 2 + 2x subscript 3 = 6
x subscript 1 + x subscript 3 = -2
2x subscript 1 + x subscript 2 + 3x subscript = 4

2. Find a basis for the null space and row space of A.
1 -1 3
A = 5 -4 -2
7 -6 4

3. A matrix in row echelon form is given. By inspection, find bases for the row and column spaces of the matrix A.

1 0 2
A = 0 0 1
0 0 0
1
Expert's answer
2020-01-08T12:31:16-0500

1. Find the vector form of the general solution of the given linear system Ax = b ; then use that result to find the vector form of the general solution of Ax = 0.

"\begin{cases}
 x_1+x_2+2x_3=6
\\x_1+x_3=-2
\\ 2x_1+x_2+3x_3=4
\end{cases}"



Let "x_3=t" , replace this in the second equation: "x_1=-2-t"

Then replace "x_3=t, x_1=-2-t" in the first and third equations:


"\begin{cases}
 (-2-t)+x_2+2t=6

\\ 2(-2-t)+x_2+3t=4
\end{cases}, \quad 
 x_2= 8-t"


The general solution of the linear system Ax=b is


"\begin{bmatrix}
x_1
\\ x_2
\\ x_3
\end{bmatrix} = 
\begin{bmatrix}
-2-t
\\ 8-t
\\t
\end{bmatrix}
= \begin{bmatrix}
-2 \\ 8 \\0 \end{bmatrix}+ t\begin{bmatrix}-1 \\ -1 \\ 0\end{bmatrix}, \quad t\in \mathbb{R}"


The general solution of the linear system Ax=0 is


"\begin{bmatrix}
x_1
\\ x_2
\\ x_3
\end{bmatrix} = 

 t\begin{bmatrix}-1 \\ -1 \\ 0\end{bmatrix}, \quad t\in \mathbb{R}"




2. Find a basis for the null space and row space of A.

"A =\begin{bmatrix} 1&-1&3
\\ 5&-4&-2
\\ 7&-6&4
\end{bmatrix}"


We reduce the matrix A as follows


"A =\begin{bmatrix} 1&-1&3
\\ 5&-4&-2
\\ 7&-6&4
\end{bmatrix} \to ^{R_2-5R_1}_{R_3-7R_1}\to

\begin{bmatrix} 1&-1&3
\\ 0&1&-17
\\ 0&1&-17
\end{bmatrix} 
\to^{R_3-R_2}_{R_1+R_2} \to 
\begin{bmatrix} 1&0&-14
\\ 0&1&-17
\\ 0&0&0
\end{bmatrix}"


The last matrix is in reduced row echelon form.


a) basis for null space

We see that the general solution of Ax=0 is

"\begin{cases} 
x_1=14x_3
\\ x_2=17x_3
\end{cases}"

where "x_3" is free variable.

The vector form solution to Ax=0 is

"\begin{bmatrix}
x_1
\\ x_2
\\ x_3
\end{bmatrix} = 

 x_3\begin{bmatrix}14 \\ 17 \\ 1\end{bmatrix}, \quad x_3\in \mathbb{R}"


The null space of the matrix A is given by

"N(A)=\{ x\in \mathbb{R}^3 | x=
 x_3\begin{bmatrix}14 \\ 17 \\ 1\end{bmatrix}, \quad x_3\in \mathbb{R}\} = Span \begin{Bmatrix} {\begin{bmatrix}14 \\ 17 \\ 1\end{bmatrix} }
\end{Bmatrix}"


We can see, that vector "\begin{bmatrix}14 \\ 17 \\ 1\end{bmatrix}" is a basis for null space.


b) basis for row space

The nonzero rows of matrix (in reduced row echelon form) is a basis for the row space of A.

Thus, "\begin{Bmatrix}
\begin{bmatrix}1 \\ 0 \\ -14\end{bmatrix},\begin{bmatrix}0 \\ 1 \\ -17\end{bmatrix}
\end{Bmatrix}" is a basis for the row space of A.




3. A matrix in row echelon form is given. By inspection, find bases for the row and column spaces of the matrix A.

"A=\begin{pmatrix} 1&0&2
\\ 0&0&1
\\ 0&0&0

\end{pmatrix}"



a) bases for the row space

The nonzero rows of a matrix in row echelon form are linearly independent.

Therefore, the row space has basis "\begin{Bmatrix}
\begin{bmatrix}1 \\ 0 \\ 2\end{bmatrix},\begin{bmatrix}0 \\ 0 \\ 1\end{bmatrix}
\end{Bmatrix}"

b) bases for the column space

The nonzero columns of a matrix in row echelon form are linearly independent.

Therefore, the column space has basis "\begin{Bmatrix}
\begin{bmatrix}1 \\ 0 \\ 0\end{bmatrix},\begin{bmatrix}2 \\ 1 \\ 0\end{bmatrix}
\end{Bmatrix}"